A 340-kg wooden raft floats on a lake. When a 80-kg man stands on the raft, it sinks 3.2 cm deeper into the water. When he steps off, the raft oscillates for a while.

Part A

What is the frequency of oscillation?

Express your answer using two significant figures.

f= ___________ Hz

Part B

What is the total energy of oscillation (ignoring damping)?

Express your answer using two significant figures.

Etotal = _____________ J
Answer #1

f = 1 / (2 pi) * sqrt (k / m)

= 1 / ( 2 pi) * sqrt ( (80*9.8/0.032) / 340)

frequency = 1.35 Hz

----------------------------------------------------

The energy of vibration E = 1/2 * k x2

= 1/2 * (80*9.8/0.032) * 0.0322

= 12.54 J